# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision

### v4

@@ 4.2 Image Recognition
Figure 4: Performance of AI models using HM, MM, and LM images for image recognition task (a) and continual learning with replay (b) and (c). (a) Performance comparison in Top-1 Accuracy between image classification models trained on HM (green), MM (yellow), and LM (blue) images is presented. The chance level is 10% for the 10-way classification (red dashed line). Error bars represent standard errors calculated over 5 independent runs. One-way ANOVA across all models yields $p<0.05$, indicating statistically significant accuracy differences among green, yellow, and blue bars. (b) Performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays, is reported as Accuracy (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for details on evaluation metrics and continual learning models. Error bars represent standard errors calculated over 5 independent runs. Two-tailed t-test analysis comparing HM vs. LM models and MM vs. LM models across all three metrics shows significant differences ($p<0.05$). (c) Accuracy at the last task (solid lines) and Forgetting (dotted lines) as a function of task numbers for continual learning models using HM (green), MM (yellow), and LM (blue) images for replays. The y-axis (Score) represents either Accuracy or Forgetting as indicated in the legend.
@@
We employ ResNet [35] as the backbone with random weights, and train it from scratch on each of the three curated subsets above over all ten classes using the Adam optimizer [44]. The initial learning rate is set to $1\times 10^{-3}$ and reduced by 50% every five epochs over a total of 20 epochs following a linear decay schedule. The resulting models are labeled as HM-model, MM-model, and LM-model based on their training subsets.
@@
We use the Cross-Entropy loss to train all three models for the 10-way classification task. Test images for object recognition are obtained from the original ImageNet [31] test set regardless of their ICMscores. As these test images lack ground truth ICMscores, we could not evaluate the top-1 accuracy of the models on individual subsets of test images stratified by ICMscores. Fig. 4(a) presents the average top-1 accuracies of the three models on all the test images for all 10 object classes over 5 random seeds.
@@
Our results in Fig. 4(a) indicate that the LM-model consistently outperforms the HM-model, achieving an average accuracy of $0.80$ compared to $0.76$ across all 10 object classes. Statistical analysis using one-way ANOVA reveal significant differences in top-1 accuracy between the two models ($p<0.05$). Surprisingly, the LM-model also surpasses the MM-model. These findings indicate that image memorability impacts image recognition, with LM images improving accuracy and HM images impairing it. We also provide the confusion matrices for the classification performances of all three models in Sec. A.3.
@@
In real-world applications, training image recognition models in computer vision benefits from using images with varying memorability. However, obtaining memorability scores for all images through human behavioral studies is often time-consuming, costly, and therefore impractical. Therefore, a computational model for predicting an image’s memorability becomes invaluable. We leverage the ICMscore predictor, introduced in Sec. 4.1, to estimate the ICMscore for all images and then use these predictions as criteria to conduct the same experiment for image recognition. From Fig. A5 (a), consistent with the findings above, recognition models trained exclusively on LM images categorized by predicted ICMscores achieve higher top-1 accuracies than the MM-model. One possible reason is that easy images for object recognition often feature prototypical representations of object classes [45], which accelerate training and foster robust, generalizable learning. As shown in Sec 3.3, easy images tend to be nearer to categorical centroids, and hence, less memorable but more beneficial for representation learning in AI models.
@@
For comparison, we also conduct the same image recognition experiment based on the classical memorability definitions from prior works [4, 5, 6, 7, 8] to divide the dataset into two subsets and utilize the predictor from [5]. The results, shown in Fig. A5 (b), indicate that the predictor proposed by [5] does not enhance image recognition accuracy. Additionally, we observe no significant difference in top-1 accuracy between models trained on the HM and LM images as predicted by [5].
@@ 4.3 Continual Learning
Continual learning, particularly class-incremental learning, is essential for AI models to adapt to new classes over time without experiencing catastrophic forgetting of previously learned classes [46]. One type of continual learning methods is based on memory replay [47, 48, 45], which involves storing a subset of previously encountered examples in a memory buffer and replaying them during subsequent tasks. Given that some images are easier to remember while others are not, we conjecture that image memorability may play a key role in improving the continual learning performance of AI models. Here, we investigate how memorability can be leveraged to select exemplar images for replay, potentially enhancing the model’s ability to retain knowledge across tasks.
@@
We structure our ICMD for continual learning by training object recognition models on the first 5 object classes in the initial task, followed by introducing 1 incremental class in each subsequent task. Since there are a total of 10 object classes, this results in 6 tasks. To examine the role of image memorability, we further divide the images from each class into HM, MM, and LM image subsets by following the same procedures in Sec. 4.2. During replay, we exclusively use images from one of these subsets, meanwhile keeping training images identical for all three continual learning models across all 6 tasks. We label the continual learning models with replays from HM, MM, and LM images as HM-models, MM-models, and LM-models respectively. We train all these 3 continual learning models in the initial task for 10 epochs, and incremental tasks for 5 epochs with an Adam optimizer using a learning rate of $1\times 10^{-3}$. The replay buffer size is set to 20 samples.
@@
Accuracy at the last task (Acc) [49] is defined as the accuracy of the model trained up to the last task $T$ and tested on all the previous tasks. The higher the Acc, the better the model performs in continual learning tasks.
Backward Transfer (BWT)[50, 51, 49] quantifies the model’s knowledge retention ability by measuring performance drop on earlier tasks after learning new ones:
@@
Forgetting (F) [49] captures the accuracy drop on the test set of the first task for the model trained at the final task. It is defined as $F=A_{1,1}-A_{T,1}$. The lower the F, the better the model performs in continual learning tasks.
Results in Fig. 4(b) show that models using LM images for replay consistently outperform those using HM and MM images in Acc, BWT and F. We also plot accuracy and forgetting as functions of task numbers for all three models in Fig. 4(c). The results indicate that the model replayed with LM images consistently exhibits reduced forgetting and achieves higher accuracy across all tasks. Together with our feature analysis in Sec. 3.3, our continual learning results align well with the previous research [52, 53] that replaying samples where their features are closer to the categorical centroids yields better continual learning performance.
Similar to the image recognition experiments in Sec. 4.2, we categorize images into HM, MM, and LM using both our ICMscore predictor and the predictor from [5]. From Fig. A5, there is no significant performance difference in replay methods using LM and HM images predicted by [5]. In contrast, our ICMscore predictor remarkably reveals significant differences, with continual learning models replayed with LM images outperforming those with HM images or models using replay images predicted by [5].
4.4 Memorability-controlled Image Editing
Figure 5: Proof-of-conept demonstration of memorability-controlled image editing. (a) During training, we fine-tune the image diffusion model, Ip2p  [12], using LoRA [13] with pairs of masked images and their corresponding ground truth images. (b) During inference, SAM [54] segments and masks the teapot area of the source image, and our fine-tuned model takes the masked image as input to edit teapot images with varying memorability. (c) The high-memorability image generator increases the ICMscore of low-memorability (LM) teapots by adding striking textures and enhancing brightness. Conversely, the LM image generator reduces the ICMscore by blurring distinctive features, making the teapots appear more ordinary with plain textures. The numbers beneath the images represent ICMscores predicted by our memorability predictor as introduced in Sec. 4.1.
Editing images with controlled memorability offers valuable applications across various research fields, such as designing advertisements for commercial products and improving students’ knowledge retention ability in education [24, 55, 56]. Next, we present a proof-of-concept demonstration of using a subset of teapot images in our ICMD to train a state-of-the-art image diffusion model capable of modifying the memorability of an original teapot image.
To adjust a teapot image’s memorability while preserving its semantic content, we train a high-memorability teapot generator (HM generator) using a subset of teapot images with high memorability (HM images) calculated from human behavior experiments (Fig. 5a). First, for each HM image, we use SAM [54] to segment and mask the teapot area with white pixels. Next, these masked images are paired with their corresponding original images to train image diffusion models. Specifically, we fine-tuned the pre-trained InstructPix2Pix model (Ip2p) [12] with LoRA [13]. The generated images by Ip2p are compared against the ground truth HM images to improve its image generation quality.
During inference, the HM generator uses “highly memorable teapot" as the text prompt and the masked teapot image as input for image editing (Fig. 5c). The masked area guides the HM generator to focus on the specified region. The generator produces highly memorable teapot images. We qualitatively evaluate the HM generator’s performance using LM teapot images from our ICMD. As illustrated in Fig. 5c, the generator effectively enhances memorability by adding visually striking textures to the original teapots. For quantitative evaluation, we use our memorability prediction model (detailed in Section 4.1) to predict ICMscores for both the original and generated images. The results indicate that the generated teapots consistently achieve higher ICMscores than the originals, demonstrating the effectiveness of our HM generator for memorability-controlled image editing.
5 Discussion
We introduce intra-class memorability, defined as the ability of individual instances within the same image class to be remembered by humans. To investigate why certain instances are more memorable than others, we contributed the ICMD dataset, comprising 5,000 images across 10 object classes. Memorability for these images is quantified using the ICMscore metric, derived from human behavior experiments in continuous object recognition tasks. To showcase the utility of ICMD, we train and evaluate AI models on subsets of the dataset, categorized by ICMscore, across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our results reveal that intra-class memorability significantly impacts feature representation learning, with low-memorability images notably improving performance in image recognition and continual learning tasks.
Building on our promising results, we outline several key directions for future research. While our study demonstrates memorability’s role in computer vision using ten object classes, expanding ICMD to hundreds of ImageNet classes would provide broader insights. Furthermore, controlling image properties—such as standardizing backgrounds—could remove confounding factors and help isolate class-specific features critical for memorability. Although our findings highlight memorability’s impact on computer vision tasks, its underlying mechanisms remain an open question. Lastly, while we focus on supervised training, future work could explore intra-class memorability in self-supervised learning frameworks.
Appendix A Appendix
A.1 Difference in Human Behavior Experiments from the Existing Literature
Previous studies typically present participants with sequences of images where each sequence contains images from multiple categories. However, image category can significantly influence memorability [4, 7, 8, 18, 19]. For example, dog images are generally more memorable than mountain images [8]. To eliminate the influence of image categories on memorability, we design a new human behavior experiment focusing on intra-class memorability. Fig. A1 compares our experimental protocol with previous approaches [4, 5, 6, 9, 7]. Unlike previous methods that mix multiple categories in a single sequence, we separate image sequences by category. Each participant views image sequences from a single category, with breaks in between image sequences from different categories to reduce category interference on memorability.
Figure A1: The differences between our experimental protocol and previous studies. Our experimental design differs from previous studies [4, 5, 6, 9, 7]. In Row 1, “Others" represents the protocol used in previous studies, where participants are exposed to images from multiple categories within a single sequence. In Row 2, “Ours" represents our protocol, which requires participants to view only images from a single category per sequence, with the dotted line indicating the break between sequences. Different colored squares represent images from distinct categories, and the numbers indicate the position of each image within its category (e.g., B2 is the second image in category B). The black "X" represents a foil image that appears only once. In our protocol, the foil image "X" belongs to the same category as the target images in the same sequence.
A.2 The Effect of Time Interval in Human Behavior Experiments
In Sec. 3.2, we state that human memory retention tends to weaken over longer intervals. That is, a longer interval $t$ makes correct identification of the repeated images more challenging. See Sec. 3.1 for the definition of interval $t$. We demonstrate this effect in Fig. A2. Fig. A2 shows that identification accuracy of the repeated images and F1 score decrease with the longer interval $t$. Additionally, we conducted an experiment with $t$ exceeding 32, during which most participants struggled to complete the task. Therefore, we empirically set $t\in\{8,16,24,32\}$ in the actual experiments.
Figure A2: Human behavior experiment scores. This figure presents human performance at different intervals of $t$ (see Sec. 3.2 for the definition of $t$). The x-axis represents $t$, and the y-axis represents the score, with the red line indicating the accuracy score and the green line indicating the F1 score [57]. The values represent the average accuracy and F1 scores [57], computed across all trials with the same $t$. The error bars represent the standard error.
A.3 Confusion Matrices in Image Recognition Experiments
Following the image recognition experiment described in Sec. 4.2, we observe that models trained on LM images perform better in image recognition tasks than HM and MM images. To further demonstrate this difference, we compare the confusion matrices of models trained on three subsets. As shown in Fig. A3, the model trained on the LM images has a lower error rate across most classes, with the largest improvement in the “bird" class, where its accuracy increases by 16% and 15% in comparison to the HM and MM models.
A.4 Performance of Models Using Memorability Predictors from the Previous Works
We replace ICMscore in ICMD with the memorability score from prior work [5] and split the dataset into two subsets (highly memorable images and low memorable images) following the same way in Sec. 4.2. In comparison, we also use our ICMscore predictor, proposed in Sec. 4.1, to divide the ICMD into three subsets LM, MM, and HM. The three subsets are then used to train image recognition models and continual learning models during replays.
Fig. A5 shows image recognition performance across the LM-model, the MM-model, and the HM-model. See Sec. 4.2 for the result discussion. In Fig. A5, we present the continual learning performance of various models with LM, MM, and HM images for replays. See Sec. 4.3 for the result discussion.

### v5

@@ 4.2 Image Recognition
We employ ResNet He et al. (2015) as the backbone with random weights, and train it from scratch on each of the three curated subsets above over all ten classes using the Adam optimizer Kingma & Ba (2017). The initial learning rate is set to $1\times 10^{-3}$ and reduced by 50% every five epochs over a total of 20 epochs following a linear decay schedule. The resulting models are labeled as HM-model, MM-model, and LM-model based on their training subsets.
@@
We use the Cross-Entropy loss to train all three models for the 10-way classification task. Test images for object recognition are obtained from the original ImageNet Deng et al. (2009) test set regardless of their ICMscores. As these test images lack ground truth ICMscores, we could not evaluate the top-1 accuracy of the models on individual subsets of test images stratified by ICMscores. Fig. 4(a) presents the average top-1 accuracies of the three models on all the test images for all 10 object classes over 5 random seeds.
Figure 4: Performance of AI models using HM, MM, and LM images for image recognition task (a) and continual learning with replay (b) and (c). (a) Top-1 accuracy of image classification models trained on HM (green), $M$ (yellow), and LM (blue) images. The chance level for 10-way classification is 10% (red dashed line). Error bars denote standard errors over 5 runs. One-way ANOVA ($p<0.05$) confirms significant accuracy differences among the three groups. (b) Continual learning performance of models replaying HM (green), $M$ (yellow), and LM (blue) images, evaluated by Accuracy (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for details. Error bars show standard errors over 5 runs. Two-tailed t-tests ($p<0.05$) reveal significant differences between LM vs. HM and LM vs. $M$ across all metrics. (c) Accuracy at the last task (solid lines) and Forgetting (dotted lines) as a function of task number for continual learning models with HM (green), $M$ (yellow), and LM (blue) replay images. The y-axis shows either Accuracy or Forgetting, as indicated in the legend.
@@
Our results in Fig. 4(a) indicate that the LM-model consistently outperforms the HM-model, achieving an average accuracy of $0.80$ compared to $0.76$ across all 10 object classes. Statistical analysis using one-way ANOVA reveal significant differences in top-1 accuracy between the two models ($p<0.05$). Surprisingly, the LM-model also surpasses the MM-model. These findings indicate that image memorability impacts image recognition, with LM images improving accuracy and HM images impairing it. We also provide the confusion matrices for the classification performances of all three models in Sec. D.1 and Fig. A4.
@@
In real-world applications, training image recognition models in computer vision benefits from using images with varying memorability. However, obtaining memorability scores for all images through human behavioral studies is often time-consuming, costly, and therefore impractical. Therefore, a computational model for predicting an image’s memorability becomes invaluable. We leverage the ICMscore predictor, introduced in Sec. 4.1, to estimate the ICMscore for all images and then use these predictions as criteria to conduct the same experiment for image recognition. See Sec.D.2 for more details. From Fig. A5 (a), consistent with the findings above, recognition models trained exclusively on LM images categorized by predicted ICMscores achieve higher top-1 accuracies than the MM-model. One possible reason is that easy images for object recognition often feature prototypical representations of object classes Singh et al. (2023), which accelerate training and foster robust, generalizable learning. As shown in Sec 3.3, easy images tend to be nearer to categorical centroids, and hence, less memorable but more beneficial for representation learning in AI models.
@@
For comparison, we also conduct the same image recognition experiment based on the classical memorability definitions from prior works Bylinskii et al. (2015); Khosla et al. (2015); Dubey et al. (2015); Goetschalckx & Wagemans (2019); Kramer et al. (2023) to divide the dataset into two subsets and utilize the predictor from Khosla et al. (2015). The results, shown in Fig. A5 (b), indicate that the predictor proposed by Khosla et al. (2015) does not enhance image recognition accuracy. Additionally, we observe no significant difference in top-1 accuracy between models trained on the HM and LM images as predicted by Khosla et al. (2015).
@@ 4.3 Continual Learning
Continual learning, particularly class-incremental learning, is essential for AI models to adapt to new classes over time without experiencing catastrophic forgetting of previously learned classes Parisi et al. (2019). One type of continual learning methods is based on memory replay Bagus & Gepperth (2021); Tee & Zhang (2023); Singh et al. (2023), which involves storing a subset of previously encountered examples in a memory buffer and replaying them during subsequent tasks. Given that some images are easier to remember while others are not, we conjecture that image memorability may play a key role in improving the continual learning performance of AI models. Here, we investigate how memorability can be leveraged to select exemplar images for replay, potentially enhancing the model’s ability to retain knowledge across tasks.
@@
We structure our ICMD for continual learning by training object recognition models on the first 5 object classes in the initial task, followed by introducing 1 incremental class in each subsequent task. Since there are 10 object classes, this results in 6 tasks. To examine the role of image memorability, we further divide the images from each class into HM, MM, and LM subsets following the procedure in Sec. 4.2. During naive replay, we exclusively use images from one of these subsets, while keeping training images identical across all three continual learning models. The models replaying HM, MM, and LM images are denoted as HM-models, MM-models, and LM-models, respectively. We train each model for 10 epochs on the initial task and 5 epochs on each incremental task using Adam with a learning rate of $1\times 10^{-3}$. The replay buffer size is fixed at 20 samples.
To benchmark the continual learning performance of these three models, we include the following three standard evaluation metrics, namely Accuracy at the last task (Acc), Backward knowledge Transfer (BWT), and Forgetfulness (F). We define $A_{t,i}$ as the accuracy of the model trained on all the tasks up to $t$ and tested on the test set of task $i$. Next, we introduce these metrics.
Accuracy at the last task (Acc) Shi et al. (2024) measures the accuracy of a model trained up to the final task $T$ and evaluated on all previous tasks; higher values indicate better continual learning. Backward Transfer (BWT) Lin et al. (2022); Díaz-Rodríguez et al. (2018); Shi et al. (2024) quantifies knowledge retention by evaluating performance changes on earlier tasks after learning new ones. Forgetting (F) Shi et al. (2024) denotes the accuracy drop on the first task when the model is trained to the last task. See Sec. E.1 for details.
Results in Fig. 4(b) show that models using LM images for replay consistently outperform those using HM and MM images in Acc, BWT and F. We also plot accuracy and forgetting as functions of task numbers for all three models in Fig. 4(c). The results indicate that the model replayed with LM images consistently exhibits reduced forgetting and achieves higher accuracy across all tasks. Together with our feature analysis in Sec. 3.3, our continual learning results align well with the previous research Rebuffi et al. (2017); Ke et al. (2024) that replaying samples where their features are closer to the categorical centroids yields better continual learning performance.
Similar to the image recognition experiments in Sec. 4.2, we categorize images into HM, MM, and LM using both our ICMscore predictor and the predictor from Khosla et al. (2015). From Fig. A6, there is no significant performance difference in replay methods using LM and HM images predicted by Khosla et al. (2015). In contrast, our ICMscore predictor remarkably reveals significant differences, with continual learning models replayed with LM images outperforming those with HM images or models using replay images predicted by Khosla et al. (2015) (see Sec. E.2).
We further conduct the above experiments on the prompt-based continual learning method Learning to Prompt (L2P) Wang et al. (2022) with an additional replay buffer. Consistent with previous results, we obtain the same conclusion that LM-based replay leads to superior performance. Full implementation details are provided in Sec. E.2.
4.4 Memorability-controlled Image Editing
Figure 5: Proof-of-conept demonstration of memorability-controlled image editing. (a) Training: the image diffusion model Ip2p Brooks et al. (2022) is fine-tuned with LoRA Hu et al. (2021) using pairs of masked and ground-truth images. (b) Inference: SAM Kirillov et al. (2023) segments and masks the teapot region of a source image, and our fine-tuned model edits the masked image to generate teapots with varying memorability. (c) The high-memorability generator raises the ICMscore of LM teapots by adding vivid textures and increasing brightness, while the LM generator lowers the ICMscore by blurring distinctive features and simplifying textures. The percentage shown on each arrow represents the relative change in memorability score, computed as the difference between the edited and original image scores, normalized by the original score. All scores are computed using our memorability regressor in (Sec. 4.1).
Editing images with controlled memorability offers valuable applications across various research fields, such as designing advertisements for commercial products and improving students’ knowledge retention ability in education Goetschalckx et al. (2019); Moran et al. (2019). We present a proof-of-concept demonstration of using a subset of teapot images in our ICMD to train an image diffusion model capable of modifying the memorability of an original teapot image.
To adjust a teapot image’s memorability while preserving its semantic content, we train a high-memorability teapot generator (HM generator) using a subset of teapot images with high memorability (HM images) calculated from human behavior experiments (Fig. 5a). First, for each HM image, we use SAM Kirillov et al. (2023) to segment and mask the teapot area with white pixels. Next, these masked images are paired with their corresponding original images to train image diffusion models. Specifically, we fine-tuned the pre-trained InstructPix2Pix model (Ip2p) Brooks et al. (2022) with LoRA Hu et al. (2021). The generated images by Ip2p are compared against the ground truth HM images to improve its image generation quality.
During inference, the HM generator uses “highly memorable teapot” as the text prompt and the masked teapot image as input for image editing (Fig. 5c). The masked area guides the HM generator to focus on the specified region. The generator produces highly memorable teapot images. We qualitatively evaluate the HM generator’s performance using LM teapot images from our ICMD. As illustrated in Fig. 5c, the generator effectively enhances memorability by adding visually striking textures to the original teapots. For quantitative evaluation, we use our memorability prediction model (detailed in Section 4.1) to predict ICMscores for both the original and generated images. The results indicate that the generated teapots consistently achieve higher ICMscores than the originals, demonstrating the effectiveness of our HM generator for memorability-controlled image editing.
5 Discussion
We introduce intra-class memorability, defined as the ability of individual instances within the same image class to be remembered by humans. To investigate why certain instances are more memorable than others, we contributed the ICMD dataset, comprising 5,000 images across 10 object classes. Memorability for these images is quantified using the ICMscore metric, derived from human behavior experiments in continuous object recognition tasks. To showcase the utility of ICMD, we train and evaluate AI models on subsets of the dataset, categorized by ICMscore, across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our results reveal that intra-class memorability significantly impacts feature representation learning, with low-memorability images notably improving performance in image recognition and continual learning tasks. Building on our promising results, we outline several key directions for future research. While our study demonstrates memorability’s role in computer vision using ten object classes, expanding ICMD to hundreds of ImageNet classes would provide broader insights. Furthermore, controlling image properties—such as standardizing backgrounds—could remove confounding factors and help isolate class-specific features critical for memorability. Although our findings highlight memorability’s impact on computer vision tasks, its underlying mechanisms remain an open question. Lastly, while we focus on supervised training, future work could explore intra-class memorability in self-supervised learning frameworks.
Appendix A Differences from Prior Work in Human Behavior Experiments
Previous studies typically present participants with sequences of images where each sequence contains images from multiple categories. However, image category can significantly influence memorability Bylinskii et al. (2015); Goetschalckx & Wagemans (2019); Kramer et al. (2023); Lu et al. (2020); Borkin et al. (2013). For example, dog images are generally more memorable than mountain images Kramer et al. (2023). To eliminate the influence of image categories on memorability, we design a new human behavior experiment focusing on intra-class memorability. Fig. A1 compares our experimental protocol with previous approaches Bylinskii et al. (2015); Khosla et al. (2015); Dubey et al. (2015); Hovhannisyan et al. (2021); Goetschalckx & Wagemans (2019). Unlike previous approaches that mix categories within a single sequence, we present image sequences one category at a time. Each participant views sequences of object instances from only one category, with breaks inserted between categories to minimize cross-category interference in memorability.
Figure A1: The differences between our experimental protocol and previous studies. Our experimental design differs from previous studies Bylinskii et al. (2015); Khosla et al. (2015); Dubey et al. (2015); Hovhannisyan et al. (2021); Goetschalckx & Wagemans (2019). In Row 1, “Others” represents the protocol used in previous studies, where participants are exposed to images from multiple categories within a single sequence. In Row 2, “Ours” represents our protocol, which requires participants to view only images from a single category per sequence, with the dotted line indicating the break between sequences. Different colored squares represent images from distinct categories, and the numbers indicate the position of each image within its category (e.g., B2 is the second image in category B). The black ”X” represents a foil image that appears only once. In our protocol, the foil image ”X” belongs to the same category as the target images in the same sequence.
Appendix B The Effect of Time Interval in Human Behavior Experiments
In Sec. 3.2, we state that human memory retention tends to weaken over longer intervals. That is, a longer interval $t$ makes correct identification of the repeated images more challenging. See Sec. 3.1 for the definition of interval $t$. We demonstrate this effect in Fig. A2. Fig. A2 shows that identification accuracy of the repeated images and F1 score decrease with the longer interval $t$. Additionally, we conducted an experiment with $t$ exceeding 32, during which most participants struggled to complete the task. Therefore, we empirically set $t\in\{8,16,24,32\}$ in real experiments.
Figure A2: Human behavior experiment scores. This figure presents human performance at different intervals of $t$ (see Sec. 3.2 for the definition of $t$). The x-axis represents $t$, and the y-axis represents the score, with the red line indicating the accuracy score and the green line indicating the F1 score Yacouby & Axman (2020). The values represent the average accuracy and F1 scores Yacouby & Axman (2020), computed across all trials with the same $t$. The error bars represent the standard error.
Appendix C Performance of Models Using ICMscore without Time Penalty
[…]
Figure A3: Performance of Models Using ICMscore without time penalty. (a) Top-1 accuracy comparison for image classification models trained on LM (blue), HM (green), and MM (yellow) images. The dashed line indicates the 20% chance level. (b) Continual learning performance showing Accuracy (left), BWT (middle), and Forgetting (right) for models using HM (green), $M$ (yellow), and LM (blue) replays. Error bars represent standard error over 5 independent runs. Evaluation metrics are detailed in Sec. 4.3.
Appendix D Image Recognition
D.1 Confusion Matrices in Image Recognition Experiments
Following the image recognition experiment described in Sec. 4.2, we observe that models trained on LM images perform better in image recognition tasks than HM and MM images. To further demonstrate this difference, we compare the confusion matrices of models trained on three subsets. As shown in Fig. A4, the model trained on the LM images has a lower error rate across most classes, with the largest improvement in the “bird” class, where its accuracy increases by 16% and 15% in comparison to the HM and MM models.
D.2 Image Recognition with Other Memorability Predictors
((a))
((b))
Figure A5: Performance comparison in Top-1 Accuracy between image classification models trained on HM, MM, and LM images using different memorability predictors. (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in Khosla et al. (2015). Within each subplot, from the top to the bottom, top-1 accuracy is reported for models trained on LM images (blue), HM images (green), and MM images (yellow). The chance is 20% for the 5-way classification (dash lines produced by the random model in red). The error bar is the standard error calculated over 5 independent runs with different random seeds.
We replace ICMscore in ICMD with the memorability score from prior work Khosla et al. (2015) and split the dataset into two subsets (highly and low memorable), following the procedure in Sec. 4.2. To examine the role of memorability in recognition, we sort images within each class by the memorability score and construct three subsets: HM (top 250), LM (bottom 250), and MM (250 randomly mixed from HM and LM). In comparison, we also apply our ICMscore predictor (Sec. 4.1) to form LM, MM, and HM subsets. These subsets are then used to train recognition models. Fig. A5 reports performance across the LM-, MM-, and HM-models; see Sec. 4.2 for discussion.
Appendix E Continual Learning
E.1 Introduction to Evaluation Metrics in Continual Learning
@@
Accuracy at the last task (Acc) Shi et al. (2024) is defined as the accuracy of the model trained up to the last task $T$ and tested on all the previous tasks. The higher the Acc, the better the model performs in continual learning tasks.
Backward Transfer (BWT)Lin et al. (2022); Díaz-Rodríguez et al. (2018); Shi et al. (2024) quantifies the model’s knowledge retention ability by measuring performance drop on earlier tasks after learning new ones:
@@
Forgetting (F) Shi et al. (2024) captures the accuracy drop on the test set of the first task for the model trained at the final task. It is defined as $F=A_{1,1}-A_{T,1}$. The lower the F, the better the model performs in continual learning.
E.2 Performance of Different Continual Learning Models